# An overview of maximal distance minimizers problem

### v2

@@ 1 Introduction
This work is devoted to solutions of the following maximal distance minimizer problem.
@@
Problem 1.1 was introduced in a very general form (also including the average distance minimization problem, which is discussed later) by Buttazzo, Oudet and Stepanov in [5]. They were motivated by urban transportation planning. The aim of an urban transportation network (for instance the set of metro or street traffic routes) is to provide the easiest access of the people to their destinations, the densities of which are supposed to be given finite Borel measures. Such a transportation network will be modeled by a closed connected set. It is reasonable to suppose that getting to a destination without using urban traffic in general would cost a single citizen an effort (estimated in time required and, at the end, also in terms of respective financial loss) proportional to the actual walking distance while the cost of using the urban traffic (the price of the ticket) is independent of the actual length of the traffic route used, and, moreover, for simplicity can be assumed to be negligible with respect to the cost of a transportation without using the urban traffic.
Later Problem 1.1 was stated in the current form by Miranda, Paolini and Stepanov in [16, 19]. Of course, we may encounter this problem in different life situations. For example, you are the mayor of a city completely occupied by rats, and you need to prepare the site for the most important international sporting event. Then you want to lay cables of a minimum length that will emit an ultrasonic rodent repeller in all areas of the city where athletes and tourists will be.
A maximal distance minimizer is a solution of Problem 1.1.
@@ 1.1 Class of problems
Maximal distance minimization problem could be considered as a particular example of shape optimization problem. A shape optimization problem is a minimization problem where the unknown variable runs over a class of domains; then every shape optimization problem can be written in the form $\min F(\Sigma):\Sigma\in A$ where $A$ is the class of admissible domains and $F()$ is the cost function that one has to minimize over $A$.
@@ 1.2 Dual problem
In a nondegenerate case (i.e. for $F_{M}(\Sigma)>0$) the primal and dual problems have the same sets of solutions for the corresponding $r$ and $l$ (see [19]) and hence an equality $F_{M}(\Sigma)=r$ is reached for a minimizer $\Sigma$.
@@ 1.4 Notation
For the both problems existence of solutions is proved easily: according to the classical Blaschke and Goł ˛ ab Theorems, the class of admissible sets is compact for the Hausdorff distance and both of the functions (maximal distance and also the average distance) is continuous for this convergence because of the uniform convergence of $x\rightarrow\mathrm{dist}\,(x,\Sigma)$.
@@
  • the Ahlfors regularity of maximum distance minimizers and, under the additional summability condition on $\phi$, the Ahlfors regularity of average distance minimizers. Gordeev and Teplitskaya [11] refine Ahlfors constants of maximum distance minimizers to the best possible, i.e. show that ${\mathcal{H}}^{1}(\Sigma\cap B_{\varepsilon}(x))=\mathrm{ord}\,_{x}\Sigma\cdot\varepsilon+o(\varepsilon)$, where $\mathrm{ord}\,_{x}\Sigma\in\{1,2,3\}$.
@@ 1.6 Local maximal distance minimizers
Any maximal distance minimizer is also a local minimizer. Usually the properties of maximal distance minimizers are also true for local maximal distance minimizers (see [11]).
@@ 2 Regularity
Figure 1: All possible variants of tangent rays at any point of a maximal distance minimizer (or blow up limits of an average distance minimizer)
It is clear that for maximal distance minimizers blow up limits also exists and are more or less the same: $\Sigma_{0}$ can be a radius, a diameter, a corner points with the angle between the segments greater or equal to $32\pi/$ 3 or a center of a regular tripod. Thus a blow up limit always belong to some plane.
@@
A key object in all the study of the average distance problem is the pull-back measure of $\mu$ with respect to the projection onto $\Sigma_{a}$, where $\Sigma_{a}$ is a solution of the average distance minimizer problem. More precisely, if $\mu$ does not charge the Ridge set (which is defined as the set of all $x\in\mathbb{R}^{d}$ for which the minimum distance to $\Sigma_{a}$ is attained at more than one point) of $\Sigma_{a}$ (this is the case for instance when $\mu$ is absolutely continuous with respect to the Lebesgue measure), then it is possible to choose a measurable selection of the projection multimap onto $\Sigma$, i.e. a map $\pi_{\Sigma}:M\rightarrow\Sigma$ such that $d(x,\Sigma)=d(x,\pi_{\Sigma_{a}})$ (this map is uniquely defined everywhere except the Ridge set). Then one can define the measure $\psi$ as being $\psi(A):=\mu(\pi_{\Sigma_{a}}^{-1}(A))$, for any Borel set $A\subset M$. In other words $\psi=\pi_{\Sigma_{a}}\sharp\mu$.
For the maximal distance minimizers in $\mathbb{R}^{d}$ we can define measure $\psi$ the similar way, but replace $M$ by $\partial B_{r}(\Sigma)$ and with $(n-1)$-dimensional Hausdorff measure as $\mu$ (or accordingly $\overline{B_{r}(\Sigma)}$ and $n$-dimensional Hausdorff measure). Thus Fig. 1 is true both for maximal and average distance minimizers.
@@ 2.3 Continuity of one-sided tangents in $\mathbb{R}^{2}$
We will say that the ray $(ax]$ is a one-sided tangent of a set $\Gamma\subset\mathbb{R}^{d}$ at a point $x\in\Gamma$ if there exists a connected component $\Gamma_{1}$ of $\Gamma\setminus\{x\}$ such that $x\in\overline{\Gamma_{1}}$ and that any sequence of points $x_{k}\in\Gamma_{1}$ with the property $x_{k}\rightarrow x$ satisfies $\angle x_{k}xa\rightarrow 0$. In this case we will also say that $(ax]$ is tangent to the connected component $\Gamma_{1}$.
@@
    For any one-sided tangent $(\bar{a}\bar{x}]$ of $\Sigma$ at $\bar{x}$ the equality $\angle((\bar{a}\bar{x}),(ax))=o_{|\bar{x}x|}(1)$ holds.
@@
    Let $(\bar{a}\bar{x}]$ be a one-sided tangent at $\bar{x}$ of any connected component of $\Sigma\setminus\{\bar{x}\}$ not containing $x$. Then $\angle((\bar{a}\bar{x}],(ax])=o_{|\bar{x}x|}(1)$.
@@ 2.5 Every $C^{1,1}$-smooth simple curve is a minimizer
For planar average distance minimizers Tilli proved in [22] that any simple $C^{1,1}$-curve is a minimizer for some given data. Paper [3] generalizes Tilli’s result on $d$-dimensional space. The same statement with a similar but much simpler explanation is true for maximal distance minimizers.
@@ 3 Explicit examples for maximal distance minimizers
A solution of Problem 3.1 is called a Steiner tree. Only vertices $x_{i}$ can have degree $1$ or $2$, all the other vertices have degree $3$ and are called Steiner points while the vertices $x_{i}$ are called terminals. Vertices of the degree $3$ are called branching points. The angle between any two adjacent edges of ${\mathcal{S}t}$ is at least $32\pi/$ 3. That means that for a branching point the angle between any two segments incident to it is exactly $32\pi/$ 3, and these three segments belong to the same 2-dimensional plane. The number of Steiner points in ${\mathcal{S}t}$ does not exceed $n-2$. A Steiner tree with exactly $22n-$ 2 vertices is called full. Every terminal point of a full Steiner tree has degree one. Proofs of the listed properties and more information on Steiner trees could be found in [10, 12].
Any maximal distance minimizer for any finite set $M\subset\mathbb{R}^{d}$ is a finite union of at most $12\sharp M-$ 1 segments. In this case maximal distance minimization problem comes down to connecting $r$-neighborhoods of all the points from $M$. If $\overline{B_{r}(a)}$ are disjoint for every $a\in M$ then a maximal distance minimizer is a Steiner tree connecting some points from $\partial B_{r}(a)$, $a\in M$.
@@ 4.1 Energetic points
where $X_{\Sigma}\subset G_{\Sigma}$ is the set of isolated energetic points (i.e. every $x\in X_{\Sigma}$ is energetic and there is a $\rho>0$ possibly depending on $x$ such that $B_{\rho}(x)\cap G_{\Sigma}=\{x\}$), $E_{\Sigma}:=G_{\Sigma}\setminus X_{\Sigma}$ is the set of non isolated energetic points and $S_{\Sigma}:=\Sigma\setminus G_{\Sigma}$ is the set of non energetic points also called the Steiner part.
Note that it is possible for a (local) minimizer in $\mathbb{R}^{d}$, $d>2$ to have no non-energetic points at all. Moreover, in some sense, any (local) minimizer does not have non-energetic points in a larger dimension:
@@
We say that a ray $(ax]$ is the energetic ray of the set $\Sigma$ with a vertex at the point $x\in\Sigma$ if there exists non stabilized sequence of energetic points $x_{k}\in G_{\Sigma}$ such that $x_{k}\rightarrow x$ and $\angle x_{k}xa\rightarrow 0$.
@@
Let $\{x_{k}\}\subset G_{\Sigma}$ and let $x\in E_{\Sigma}$ be the limit point of $\{x_{k}\}$: $x_{k}\rightarrow x$. By basic property of energetic points for every point $x_{k}\in G_{\Sigma}$ there exists a point $y_{k}\in M$ (may be not unique) such that $\mathrm{dist}\,(x_{k},y_{k})=r$ and $B_{r}(y_{k})\cap\Sigma=\emptyset$. In this case we will say that $y_{k}$ corresponds to $x_{k}$.
@@ 4.2 Derivation in the picture
During this subsection $M$ is a planar convex closed smooth curve with the radius of curvature greater than $r$.
@@
    Let $x$ be an isolated energetic point of degree 1 (i.e. $x$ is the end of the segment $[xz]\subset\Sigma$) with unique $y(x)$. Then $x$, $z$ and $y(x)$ lie on the same line.
@@
    Let $x$ be an isolated energetic point of degree 2 (i.e. $x$ is the end of the segments $[xz_{1}]$ and $[xz_{2}]\subset\Sigma$) with unique $y(x)$. Then $\angle z_{1}xy(x)=\angle y(x)xz_{2}$.
@@
Let $y\in M$ be a point such that $B_{r}(y)\cap\Sigma=\emptyset$ and $\partial{B_{r}(y)}$ contains energetic points $x_{1}$ and $x_{2}$. Define $n_{y}=\partial B_{r}(y)\cap M_{r}$. Then
@@
    derivatives of length of $\Sigma$ in neighborhoods of $x_{1}$ and $x_{2}$ when moving $y$ along $M$ are equal.
In [8] the derivative of length of $\Sigma$ in a neighborhood of $x$ when moving $y$ along $M$ is calculated. The derivative depends on the behavior of $\Sigma$ in the neighborhood of $x$. Since $M$ has a large radius of curvature, $\partial B_{r}(x)$ intersects $M$ in at most 2 points, so every energetic point has at most 2 corresponding $y$. Thus, we have the following 4 cases.
1. $x$ has order 1, and there is unique corresponding $y(x)$.
Then the derivative is equal to
$\cos\alpha,$
where $\alpha=\angle([xy(x)),l)$, $l$ is a tangent ray to $M$ at point $y(x)$, in the direction of increasing $\gamma(x)$.
2. $x$ has order 2, and the unique corresponding $y(x)$.
Since $x$ has order 2, $B_{\varepsilon}(x)\cap\Sigma=[xz_{1}]\cup[xz_{2}]$ for small enough $\varepsilon>0$. Then the derivative is equal to
$2\cos\alpha\cos\frac{\angle z_{1}xz_{2}}{2},$
where $\alpha=\angle([xy(x)),l)$, $l$ is a tangent ray to $M$ at point $y(x)$, in the direction of increasing $\gamma(x)$.
3. The degree of point $x$ is 1, and there are two distinct points $y_{1}(x)$ and $y_{2}(x)$.
Define $\alpha=\angle xy_{1}(x)y_{2}(x)=\angle xy_{2}(x)y_{1}(x)$ and let $\delta$ be the angle between $y_{1}(x)y_{2}(x)$ and $M$.
Let $\beta$ be the angle between $[zx]$ and the $x$ axis. Then the derivative is equal to
$\frac{\cos(\alpha+\delta)\sin(\alpha+\beta)}{\sin(2\alpha)}.$
4. The degree of point $x$ is 2, and there are two distinct points $y_{1}(x)$ and $y_{2}(x)$.
The derivative is equal to
$\frac{\cos(\alpha+\delta)}{\sin(2\alpha)}(\sin(\alpha+\beta)+\sin(\alpha+\gamma)),$
where $\beta$ and $\gamma$ are the angles between the $x$ axis and the segments $[z_{1}x]$ and $[z_{2}x]$, respectively, $\alpha$ and $\delta$ are similar to the previous case.

### v3

@@ 1 Introduction
This work is devoted to solutions to the following maximal distance minimizer problem.
@@
Problem 1.1 with the average distance minimization problem (which is discussed later) were introduced in a very general form by Buttazzo, Oudet and Stepanov in [5]. They were motivated by urban transportation planning. The aim of an urban transportation network (for instance the set of metro or street traffic routes) is to provide the easiest access of the people to their destinations, the densities of which are supposed to be given finite Borel measures. Such a transportation network will be modeled by a one-dimensional closed connected set. It is reasonable to suppose that getting to a destination without using urban traffic in general would cost a single citizen an effort (estimated in time required and, at the end, also in terms of respective financial loss) proportional to the actual walking distance while the cost of using the urban traffic (the price of the ticket) is independent of the actual length of the traffic route used, and, moreover, for simplicity can be assumed to be negligible with respect to the cost of a transportation without using the urban traffic.
Later Problem 1.1 was stated in the current form by Miranda, Paolini and Stepanov in [16, 20]. Of course, we may encounter this problem (and its dual form) in different life situations. For example, you are the mayor of a city completely occupied by rats, and you need to prepare the site for the most important international sporting event. Then you want to lay cables of a minimum length that will emit an ultrasonic rodent repeller in all areas of the city where athletes and tourists will be.
A maximal distance minimizer is a solution to Problem 1.1.
@@ 1.1 Class of problems
Maximal distance minimization problem could be considered as a particular example of shape optimization problems. A shape optimization problem is a minimization problem where the unknown variable runs over a class of domains; then every shape optimization problem can be written in the form $\min F(\Sigma):\Sigma\in A$ where $A$ is the class of admissible domains and $F()$ is the cost function that one has to minimize over $A$.
@@ 1.2 Dual problem
In a nondegenerate case (i.e. for ${\mathcal{H}}^{1}(\Sigma)>0$) the primal and dual problems have the same sets of solutions for the corresponding $r$ and $l$ (see [20]) and hence an equality $F_{M}(\Sigma)=r$ is reached for a solution $\Sigma$ to Problem 1.2.
@@ 1.3 The first parallels with the average distance minimization problem
Maximal distance minimization problem is somehow similar to another shape optimization problem: average distance minimization problem (see the survey of Lemenant [14]) and it seems interesting to compare the known results and open questions concerning these two problems. In the average distance minimization problem’s statement the admissible set $A$ is the same as in maximal distance minimization problem, but the cost function is defined as
@@
where $A:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ is a nondecreasing function and $\phi()$ is a finite nonnegative measure with compact nonempty support in $\mathbb{R}^{d}$.
@@ 1.4 Notation
A regular tripod is a union of three segments $[ax]\cup[bx]\cup[cx]$ with $\angle axb=\angle axc=\angle bxc=2\pi/3$. Note that a regular tripod is always coplanar.
@@
For the both problems the existence of solutions can be obtained in a straightforward way: according to the classical Blaschke and Goł ˛ ab theorems, the class of admissible sets is compact for the Hausdorff distance and both of the functions (maximal distance and also the average distance) is continuous for this convergence because of the uniform convergence of $x\rightarrow\mathrm{dist}\,(x,\Sigma)$.
@@
  • the Ahlfors regularity of maximum distance minimizers (in the plane the constants are absolute) and, under the additional summability condition on $\phi$, the Ahlfors regularity of average distance minimizers. Gordeev and Teplitskaya [11] refine Ahlfors constants of planar maximum distance minimizers to the best possible, i.e. show that ${\mathcal{H}}^{1}(\Sigma\cap B_{\varepsilon}(x))=\mathrm{ord}\,_{x}\Sigma\cdot\varepsilon+o(\varepsilon)$, where $\mathrm{ord}\,_{x}\Sigma\in\{1,2,3\}$. Also, Gordeev and Teplitskaya proved a ‘‘local’’ analogue in $\mathbb{R}^{d}$, namely the same result with $\varepsilon$ depending on $x$.
@@
1.6 Complexity of the problem
The corresponding discretization of Problem 1.1 is NP-complete (see Problem 5.3 and Section 5.3), this means that Problem 1.1 itself is NP-hard.
Even in the plane and for a general 3-point set $M=\{a,b,c\}$ a solution to Problem 1.1 cannot be constructed by a straightedge (ruler) and a compass. The reason is that for a large enough $\angle abc$ and a proper $r$ Problem 1.1 coincides with Alhazen’s billiard problem (a reflection from a spherical mirror), for which the unconstructability was shown by Neumann [18]. This differs the situation from the Steiner tree problem (Problem 3.1), see [15].
@@ 1.7 Local maximal distance minimizers
Any maximal distance minimizer is also a local minimizer. Usually the regularity properties of maximal distance minimizers are also true for local maximal distance minimizers (see [11]).
@@ 2 Regularity
Figure 1: All possible variants of tangent rays at any point of a maximal distance minimizer or blow up limits of an average distance minimizer
It is clear that for maximal distance minimizers blow up limits also exists and are more or less the same: $\Sigma_{0}$ can be a radius, a diameter, a union of two segments with the angle between the segments greater or equal to $32\pi/$ 3 or a center of a regular tripod. Thus a blow up limit always belongs to some 2-dimensional plane.
@@
A key object in all the study of the average distance problem is the pull-back measure of $\mu$ with respect to the projection onto $\Sigma_{a}$, where $\Sigma_{a}$ is a solution to the average distance minimizer problem. More precisely, if $\mu$ does not charge the Ridge set (which is defined as the set of all $x\in\mathbb{R}^{d}$ for which the minimum distance to $\Sigma_{a}$ is attained at more than one point) of $\Sigma_{a}$ (this is the case for instance when $\mu$ is absolutely continuous with respect to the Lebesgue measure), then it is possible to choose a measurable selection of the projection multimap onto $\Sigma$, i.e. a map $\pi_{\Sigma}:M\rightarrow\Sigma$ such that $d(x,\Sigma)=d(x,\pi_{\Sigma_{a}})$ (this map is uniquely defined everywhere except the Ridge set). Then one can define the measure $\psi$ as being $\psi(A):=\mu(\pi_{\Sigma_{a}}^{-1}(A))$, for any Borel set $A\subset M$. In other words $\psi=\pi_{\Sigma_{a}}\sharp\mu$.
For the maximal distance minimizers in $\mathbb{R}^{d}$ we can define measure $\psi$ in the similar way, but replace $M$ by $\partial B_{r}(\Sigma)$ and with $(n-1)$-dimensional Hausdorff measure as $\mu$ (or accordingly $\overline{B_{r}(\Sigma)}$ and $n$-dimensional Hausdorff measure). Thus Fig. 1 works both for maximal and average distance minimizers.
@@ 2.3 Continuity of one-sided tangents in $\mathbb{R}^{2}$
We say that the ray $(ax]$ is a one-sided tangent of a set $\Gamma\subset\mathbb{R}^{d}$ at a point $x\in\Gamma$ if there exists a connected component $\Gamma_{1}$ of $\Gamma\setminus\{x\}$ such that $x\in\overline{\Gamma_{1}}$ and any sequence of points $x_{k}\in\Gamma_{1}$ with the property $x_{k}\rightarrow x$ satisfies $\angle x_{k}xa\rightarrow 0$. In this case we also say that $(ax]$ is tangent to the connected component $\Gamma_{1}$.
@@
    For any one-sided tangent $(\bar{a}\bar{x}]$ of $\Sigma_{1}$ at $\bar{x}$ the equality $\angle((\bar{a}\bar{x}),(ax))=o_{|\bar{x}x|}(1)$ holds.
@@
    Let $(\bar{a}\bar{x}]$ be a one-sided tangent at $\bar{x}$ of any connected component of $\Sigma_{1}\setminus\{\bar{x}\}$ not containing $x$. Then $\angle((\bar{a}\bar{x}],(ax])=o_{|\bar{x}x|}(1)$.
@@ 2.5 Every $C^{1,1}$-smooth simple curve is a minimizer
For planar average distance minimizers Tilli proved in [23] that any simple $C^{1,1}$-curve is a minimizer for a proper input. Paper [3] generalizes Tilli’s result on $d$-dimensional space. The same statement with a similar but much simpler explanation is true for maximal distance minimizers.
@@ 3 Explicit examples for maximal distance minimizers
A solution ${\mathcal{S}t}$ to Problem 3.1 is called a Steiner tree. It can be represented as a plane graph, such that its set of vertices contains $P$, and all its edges are line segments. This graph is connected and does not contain cycles, i.e. is a tree, which explains the naming of $S$. It is known that the maximum vertex degree of this graph is 3. Only vertices $x_{i}$ can have degree $1$ or $2$, all the other vertices have degree $3$ and are called Steiner points while the vertices $x_{i}$ are called terminals. Vertices of the degree $3$ are called branching points. The angle between any two adjacent edges of ${\mathcal{S}t}$ is at least $32\pi/$ 3. That means that for a branching point the angle between any two segments incident to it is exactly $32\pi/$ 3, and these three segments belong to the same 2-dimensional plane. The number of Steiner points in ${\mathcal{S}t}$ does not exceed $n-2$. A Steiner tree with exactly $22n-$ 2 vertices is called full. Every terminal point of a full Steiner tree has degree one. Full trees are indecomposable in the sense that a full tree cannot be represented as a union of Steiner trees for non-empty $P_{1}$ and $P_{2}$ such that $P_{1}\cup P_{2}=P$.
We define a locally minimal tree for $P$ as a connected compact acyclic set $S$, which contains $P$, and such that for any $x\in S\setminus P$ there is a neighborhood $U\ni x$ such that $S\cap U$ coincides with the Steiner tree on the set of points $(S\cap\partial U)\cup(P\cap U)$. A locally minimal tree retains the following properties of a Steiner tree: it is the union of a finite set of segments; the angle between any two adjacent segments is at least $32\pi/$ 3.
Proofs of the listed properties and more information on Steiner trees could be found in [10, 12].
Any maximal distance minimizer for any finite set $M\subset\mathbb{R}^{d}$ is a finite union of at most $32\sharp M-$ 3 segments. In this case maximal distance minimization problem comes down to connecting $r$-neighborhoods of all the points from $M$. If $\overline{B_{r}(a)}$ are disjoint for every $a\in M$ then a maximal distance minimizer is a Steiner tree connecting some points from $\partial B_{r}(a)$, $a\in M$.
@@ 4.1 Energetic points
where $X_{\Sigma}\subset G_{\Sigma}$ is the set of isolated energetic points (i.e. every $x\in X_{\Sigma}$ is energetic and there is a $\rho>0$ possibly depending on $x$ such that $B_{\rho}(x)\cap G_{\Sigma}=\{x\}$), $E_{\Sigma}:=G_{\Sigma}\setminus X_{\Sigma}$ is the set of non-isolated energetic points and $S_{\Sigma}:=\Sigma\setminus G_{\Sigma}$ is the set of non-energetic points also called the Steiner part.
Note that it is possible for a (local) minimizer in $\mathbb{R}^{d}$, $d>2$ to have no non-energetic points at all. Moreover, in some sense, any (local) minimizer does not have non-energetic points in an embedding into a larger dimension:
@@
We say that a ray $(ax]$ is the energetic ray of the set $\Sigma$ with a vertex at the point $x\in\Sigma$ if there exists a sequence of energetic points $x_{k}\in G_{\Sigma}$ such that $x_{k}\rightarrow x$ and $\angle x_{k}xa\rightarrow 0$.
@@
Let $\{x_{k}\}\subset G_{\Sigma}$ and let $x\in E_{\Sigma}$ be the limit point of $\{x_{k}\}$: $x_{k}\rightarrow x$. By basic property of energetic points for every point $x_{k}\in G_{\Sigma}$ there exists a point $y_{k}\in M$ (may be not unique) such that $\mathrm{dist}\,(x_{k},y_{k})=r$ and $B_{r}(y_{k})\cap\Sigma=\emptyset$. Recall that it is called that $y_{k}$ corresponds to $x_{k}$.
@@ 4.2 Derivation in the picture
During this subsection $M$ is a planar convex closed smooth curve with the radius of curvature greater than $r$; we follow paper [8]. Assume that the left and right neighborhoods of $y\in M$ are contained in $r$-neighborhoods of different energetic points $x_{1}$, $x_{2}\in\Sigma$. We write conditions on the behavior of $\Sigma$ in the neighborhoods of $x_{1}$ and $x_{2}$ under the assumption by moving $y$ along $M$, which follow from the optimality of $\Sigma$.
We start with a preliminary lemma.
@@
    Assume that $x$ is an isolated energetic point of degree 1 (i.e. $x$ is the end of the segment $[xz]\subset\Sigma$) with a unique corresponding point $y(x)\in M$. Then $x$, $z$ and $y(x)$ lie on the same line.
@@
    Assume that $x$ is an isolated energetic point of degree 2 (i.e. $x$ is the end of distinct segments $[xz_{1}]$ and $[xz_{2}]\subset\Sigma$) with a unique corresponding $y(x)$. Then $\angle z_{1}xy(x)=\angle y(x)xz_{2}$.
Recall that $M_{r}$ denotes the inner part of the boundary of $B_{r}(M)$. Let $S$ be the closure of a connected component of $\Sigma\setminus\operatorname{conv}M_{r}$. For a connected $\sigma\subset\Sigma$ define the part of $M$ which is covered by $\sigma$ as $Q(x)=\overline{B_{r}(x)}\cap M$; by the restrictions on $M$ the set $Q(x)$ is always an arc of $M$. We denote one of the ends of the arc $Q(S):=\overline{B_{r}(S)}\cap M$ by $y_{1}$. Let $x\in\left(\partial B_{r}(y_{1})\setminus\operatorname{conv}M_{r}\right)\cap S$ be the energetic point for which $y_{1}$ is corresponding. Clearly $S$ is a locally minimal tree for its energetic points and points from $S\cap M_{r}$. Also it is clear that $x$ cannot have more than two corresponding points. If there are two corresponding points, we denote the second by $y_{2}$. Let us denote the degree of $x$ by $d\in\{1,2\}$, and the number of points corresponding to $x$ by $k\in\{1,2\}$. Thus, there are 4 possible cases, each of which we will consider in detail below.
Let us fix a $l>0$ such that $\overline{B_{l}(x)}\cap\Sigma$ is the union of $d$ segments of the form $[z_{i}x]$, $z_{i}\in\partial B_{l}(x)$, $1\leq i\leq d$. For a sufficiently small $0\leq\varepsilon<\varepsilon_{0}(l,r,\{y_{j}\}_{j=1}^{k},\{z_{i}\}_{i=1}^{d},x)$ denote by $y_{1}^{\varepsilon}$ the point obtained by shifting the point $y_{1}$ along $M$ by $\varepsilon$ (that is, such that the arc $M$ with ends at $y_{1}$ and $y_{1}^{\varepsilon}$ has length $\varepsilon$) in such a direction that $y_{1}^{\varepsilon}\not\in Q(S)$. For a sufficiently small in modulus $0>\varepsilon>-\varepsilon_{0}(l,r,\{y_{j}\}_{j=1}^{k}\{z_{i}\}_{i=1}^{d},x)$ we denote by $y_{1}^{\varepsilon}$ the point obtained by shifting $y_{1}$ along $M$ by $-\varepsilon$ in the opposite direction (that is, in such a way that $y_{1}^{\varepsilon}\in Q(S)$). Let us denote $y_{1}^{0}=y_{1}$. In the case of $k=2$, we denote $y_{2}^{\varepsilon}=y_{2}$ for any $\varepsilon$. Put
$\Gamma(\varepsilon)=\min_{x^{\prime}}\sum_{i=1}^{d}|z_{i}x^{\prime}|,$
where the minimum is taken over all points $x^{\prime}$ such that $|y_{j}^{\varepsilon}x^{\prime}|=r$ for every $1\leq j\leq k$. Let us denote by $x_{\varepsilon}$ a point at which the value $\Gamma(\varepsilon)$ is reached.
Note that $x_{0}=x$, since $\Sigma$ is a minimizer. The derivative $\Gamma(\varepsilon)$ at the origin $\Gamma^{\prime}(0)$ will be called the derivative of the length $\Sigma$ in the neighborhood of the point $x$ as $y_{1}^{\varepsilon}$ moves along $M$. We calculate this derivative in each of the four cases.
@@
Let $y\in M$ be a point such that $B_{r}(y)\cap\Sigma=\emptyset$, and $\partial{B_{r}(y)}$ contains energetic points $x_{1}$ and $x_{2}$. Define $n_{y}=\partial B_{r}(y)\cap M_{r}$. Then
@@
    derivatives of the length of $\Sigma$ in neighborhoods of $x_{1}$ and $x_{2}$ when moving $y$ along $M$ are equal.
All the angles defined in the following cases are assumed to belong to $[0,\pi/2]$.
Case 1: $d=1$, $k=1$.
Let the angle between $(xy_{1}]$ and $M$ be equal to $\alpha$. Then $\Gamma^{\prime}(0)=\cos\alpha$.
Case 2: $d=2$, $k=1$.
Let the angle between $(xy_{1}]$ and $M$ be equal to $\alpha$ and let $\beta=\frac{1}{2}\angle z_{1}xz_{2}$. Then $\Gamma^{\prime}(0)=2\cos\alpha\cos\beta$.
For the next two cases we need more notation. Triangle $xy_{1}y_{2}$ is isosceles, denote $\angle xy_{1}y_{2}=\angle xy_{2}y_{1}=:\alpha$. Let us introduce the following coordinates: midpoint $o$ of the segment $[y_{1}y_{2}]$ is the origin of coordinates; $X$ axis is aligned with the beam $[y_{2}y_{1})$; the $Y$ axis is codirected with the ray $[ox)$. In particular, we have
$o=(0,0),\quad x=(0,r\sin\alpha),\quad y_{1}=(r\cos\alpha,0),\quad y_{2}=(-r\cos\alpha,0).$
Let the angle between straight line $(y_{1}y_{2})$ and $M$ at point $y_{1}$ be equal to $\delta$.
Case 3: $d=1$, $k=2$.
Let us denote by $\beta$ the angle between $[z_{1}x]$ and the $X$ axis. Then
$\Gamma^{\prime}(0)=\frac{\cos(\alpha+\delta)\sin(\alpha+\beta)}{\sin(2\alpha)}.$
Case 4: $d=2$, $k=2$.
As in case 3, let $\beta$ denote the angle between $[z_{1}x]$ and the $X$ axis; similarly, we denote by $\gamma$ the angle between $[z_{2}x]$ and the $X$ axis. Then
$\Gamma^{\prime}(0)=\frac{\cos(\alpha+\delta)}{\sin(2\alpha)}(\sin(\alpha+\beta)+\sin(\alpha+\gamma)).$